OpenStudy (anonymous):

Given that f(x) = 5x − 10 and g(x) = x + 3, solve for f(g(x)) when x = −1.

OpenStudy (alekos):

just substitute g(x) = x+3 in f(g(x)) and go from there........

OpenStudy (anonymous):

what is g(-1)

OpenStudy (anonymous):

or you can do like this. f(g(x))=f(x+3)=5(x+3)-10 simplify and then find the solution at x=-1